I am trying to figure out how to solve this problem using the substitution method. I know that it equals 1000, but I can't figure out how to show the proper work. Any help of guidance would be appreciated. Thank you.
60+.20(x-550)=120+.15(x-800)

Answer by ewatrrr(24785)

 
Hi
solving by:
60 + .20(x-550)=120 +.15(x-800)
1) removing the parenthesis
60+.20x-110 =120+.15x-120
2)combining Like terms
.20x - 50 = .15x
3) Isolating the variable on one side of the EQ
.05x = 50
x = 50/.05
x = 1000